OpenStudy (anonymous):

Really need help! If a line goes through points (-2, 9) and (0, 3), what is its slope? -3 3 6 -1/3

OpenStudy (mathstudent55):

Slope = difference in y divided by difference in x. Subtract the y-coordinates. In the same order, subtract the x-coordinates. Divide the first result by the second result. That is the slope.

OpenStudy (anonymous):

So 9-3=6 -2-0=-2 So 6 divided by 2? =3? or -3?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

You are correct in the operations you did. 6/(-2) = -3 The slope is -3. When you divide positive by negative, you get negative.

OpenStudy (anonymous):

Thanks!

OpenStudy (mathstudent55):

You're welcome.